OpenStudy (anonymous):

Find x is y = -2 4x-6/y=15

OpenStudy (anonymous):

Sub in -2 for y.

OpenStudy (anonymous):

4x - 6 / -2

OpenStudy (anonymous):

2x + 3 = 15

OpenStudy (anonymous):

2x = 12

OpenStudy (anonymous):

x = 6

OpenStudy (anonymous):

4x+3=15 =>4x=15-3=12 =>x=12/4=3 so x=3

OpenStudy (anonymous):

Zai you forget the / -2

OpenStudy (anonymous):

i don't forget dear..review your answer and question carefullty too...

OpenStudy (anonymous):

Oh, and I messed up. -2x + 3 = 15 -2x = 12 x = -6

OpenStudy (anonymous):

sory its +6 not -6..you are again missing...

OpenStudy (anonymous):

4x / -2 = -2x

OpenStudy (anonymous):

-2x = 12 x = -6

OpenStudy (anonymous):

Thank you guys so much . >.<

OpenStudy (anonymous):

Trust me, it's -6.

OpenStudy (anonymous):

-6 isn't an answer though..

OpenStudy (anonymous):

its 6 not -...coz you have to divide nd term which has y in denominator...not st term...!!!